OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -7, an = 2 • an-1

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

i got -7, -14, -28, -56, -112, -224

OpenStudy (anonymous):

Let me check if you are right or not...

OpenStudy (solomonzelman):

Wrong way, will you find the 1000th term the same way ?

OpenStudy (anonymous):

What is the condition given, write it properly..

OpenStudy (anonymous):

oh!!!

OpenStudy (anonymous):

-7, -14, -12, -10, -8, -6

OpenStudy (anonymous):

now it makes sense

OpenStudy (solomonzelman):

\(\large\color{black}{ a_n=a_1 \times r^{n-1}}\) \(\large\color{black}{ a_6=a_1 \times r^{6-1}}\) \(\large\color{black}{ a_6=(-7) \times 2^{5}=?}\)

OpenStudy (anonymous):

\[a_n = 2 \cdot a_{n-1}\] Is it??

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

I showed how to find the 6th term -:(

OpenStudy (solomonzelman):

What water in eyes said, is just a recursive way of writing that the common ratio is 2.

OpenStudy (anonymous):

Hey, is it GP or AP??

OpenStudy (anonymous):

-224

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

Yeah, nice.. :)

OpenStudy (anonymous):

wait so i was right before? now im confused

OpenStudy (solomonzelman):

Can you find the 10th term in this sequence ?

OpenStudy (solomonzelman):

You weren't

OpenStudy (solomonzelman):

You can't just keep multiplying like this, the teacher will give you a zero for that.

OpenStudy (anonymous):

-7, -14, -28, -56, -112, -224 it says first 6 terms!

OpenStudy (solomonzelman):

Yes, but this is not the right approach though.

OpenStudy (anonymous):

wait so why are we doing the 10th then?

OpenStudy (solomonzelman):